OpenStudy (georgia412):

a line has a gradient of 6 passes through the points with coordinates (-1,6) and (7,a). find the value of a? How do you work out what a is?

OpenStudy (triciaal):

one approach the gradient is the same as the slope slope = change in y / corresponding change in x change in y = (y2 - y1) / change in x = (x2 - x1) for any 2 points on the line (x1, y1) and (x2, y2) for this problem ,